OpenStudy (sonyalee77):

Solve the inequality and then graph its solution: v+3<=2(v-7)

OpenStudy (sonyalee77):

@texaschic101 @wolfe8 @eliassaab

OpenStudy (sonyalee77):

@JamesWolf

OpenStudy (anonymous):

\[ v+3 \le 2 v -14\\ 3+14 \le 2v -v\\ \] Can you finish it?

OpenStudy (anonymous):

\[ 17 \le v \]

OpenStudy (sonyalee77):

v=17?

OpenStudy (sonyalee77):

I mean V<=17?

OpenStudy (sonyalee77):

@eliassaab

OpenStudy (anonymous):

No \[ v \ge 17 \]

OpenStudy (sonyalee77):

v>=17?

OpenStudy (sonyalee77):

Sorry your equation tht u made won't pop up i think tht there is something wrong with openstudy.

OpenStudy (anonymous):

v >= 17

OpenStudy (anonymous):

v greater or equal to 17